If a circuit has 50 Hz frequency and its peak value is 20V. Calculate the sinusoidal voltage. Also calculate the time for one cycle and draw the corresponding voltage versus time graph.



1
Expert's answer
2020-08-21T11:00:22-0400

Given data

frequency(f)=50 Hz

peak value of voltage(V0)=20 Volt

Angular frequency of sinusoidal voltage wave

"\omega =2 \pi f"

"\implies \omega=2\pi\times50=100\pi" rad/sec


standard sinusoidal form of voltage is given by


"V=V_{0}\sin\omega t"

putting value of V0 and "\omega" then

"{V=20\sin100\pi t}"


time period for one cycle is given by


"T=\frac{1}{f}= \frac{2\pi}{\omega}"


Now "T=\frac{1}{50}=0.02 sec"

therefore sinusoidal voltage form is "V=V_{0}\sin100\pi t" and time period for one cycle "T=0.02 sec".

"\implies" voltage versus time graph is below